OpenStudy (anonymous):

Why is 7x-8y/8y-7x=-1. Could someone please explain to me??

OpenStudy (anonymous):

7x-8y/8y-7x=-1 (multiply both sides by the denominator) 7x-8y=-1(8y-7x) 7x-8y=-8y+7x or 7x-8y=7x-8y

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

yw